OpenStudy (anonymous):

need help on a problem. 3x + 2y = 6 3x + 3=y these are my choice for this problem.. need quick answer. (-1,-6) (0,3) (1,6) (3,3)

HanAkoSolo (jamierox4ev3r):

I have to give you a process, you know, code of conduct 3x+2(3x+3)=6 3x+6x+6=6 9x+6=6 -6 -6 9x=0 x=0 3(0)+3=y 0+3=y y=3

HanAkoSolo (jamierox4ev3r):

so if x=0 and y=3, what are your final points? @PeytonPerez372 ?

OpenStudy (anonymous):

my final points would be 3,3??

HanAkoSolo (jamierox4ev3r):

no, your final point would be (0,3) :)

OpenStudy (anonymous):

ohhh... wow :O I fell dumb...